What is the probability that the random variable is less than or equal to 75 if the mean is 20 and the standard deviation is 4?
Answer by stanbon(75887)

What is the probability that the random variable is less than or equal to 75 if the mean is 20 and the standard deviation is 4?
-------------------
Find the z-value of 75:
z(75) = (75-20)/4 = 55/4 = 13.75
Note: That means 75 is 13.75 standard deviations above the mean.
----
P(x<= 75) = P(z <= 13.75) = 1 on my calculator
===================================
Cheers,
Stan H.